OpenStudy (anonymous):

If f is positive and concave upward on an interval I, what is the concavity of the function g(x) = [f(x)]^2 on the interval I? Justify your claim. I am having troubles doing this question if it is squared wouldn't it all be concave down? Thanks.

OpenStudy (misty1212):

\[f(x)=x^2\] fits that bill

OpenStudy (misty1212):

and \[f^2(x)=x^4\] is concave up for sure

OpenStudy (anonymous):

\[g(x)=[f(x)]^2\implies g'(x)=2f(x)f'(x)\implies g''(x)=2[f'(x)]^2+2f(x)f''(x)\] Given that \(f\) is positive and concave up in some interval \(I\), you know that \(f(x)>0\) and \(f''(x)>0\). What sign must \(g''(x)\) have?

OpenStudy (anonymous):

It must be positive?

OpenStudy (anonymous):

Right, so if the second derivative is positive, what do you know about the concavity of \(g\)?

OpenStudy (anonymous):

It is concave up!

OpenStudy (anonymous):

Precisely

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

yw